OpenStudy (gdgutierrez):

Please help. I have a question like this on a upcoming midterm and I can't seem to find the answer. Given Polar equation of curve C. C: r= sinθ 0<= θ <= pi Find the point, on C such that: Tangent line: a)Horizontal b)Vertical

OpenStudy (gdgutierrez):

so far I have x= sinθcosθ, y= sinθsinθ-> sin^2(θ) dx = cos(2θ) dy = sin(2θ) \[\frac{ dy }{ dx} =\frac{ dy/dθ }{ dx/dθ}=\frac{ \sin(2θ) }{ \cos(2θ) }\] a) horizontal : sin(2θ) = 0

OpenStudy (nincompoop):

tan?

OpenStudy (gdgutierrez):

I don't seem to understand what you mean by "tan"

OpenStudy (nincompoop):

sin/cos = tan

OpenStudy (gdgutierrez):

I don't think it can work that way.

OpenStudy (nincompoop):

k i just woke up so bear with me

OpenStudy (gdgutierrez):

It's cool, no problem. I say it won't because I need to find the horizontal and vertical lines

OpenStudy (nincompoop):

I feel like your question changed

OpenStudy (nincompoop):

nvm

OpenStudy (gdgutierrez):

How about L'hopitals?

OpenStudy (perl):

we know that x = r cos theta, but r is a function of theta itself x = r(theta) * cos theta y = r(theta) * sin theta . ok im going to use t instead of theta. x = r(t) * cos(t) dx/dt = r'(t) * cos(t) + r(t) * (-sin(t)) dy/dt = r'(t)*sin(t) + r(t) * cos(t)

OpenStudy (gdgutierrez):

So product rule...ooooooh figures... Thanks!

OpenStudy (perl):

but your approach is interesting too

OpenStudy (gdgutierrez):

Yours makes more sense haha. Thanks for the link!

OpenStudy (nincompoop):

i am trying to figure out what she was doing and looked it up but can't make sense

OpenStudy (nincompoop):

she's turning it into polar coordinates is what I gather

OpenStudy (perl):

she did, since r= sin theta (that is given in the directions.) then the conversion equation is x = r cos theta, ---> x = sin theta * cos theta

OpenStudy (perl):

then identity: sin (2*theta) = 2 sin theta cos theta so it should be x = sin (2theta) / 2

OpenStudy (nincompoop):

k now I see

OpenStudy (perl):

so dx/dtheta = cos (2theta). that part is correct

OpenStudy (perl):

and dy/dtheta = 2 sin(theta) * cos(theta) = sin (2theta) . that is correct too

OpenStudy (perl):

now you could say , since tan x = sin x / cos x then dy/dx = tan (2theta) . but its easier if we keep it as sin(2theta)/cos(2theta) since then you can see that horizontal tangents occur when the numerator = 0, vertical tangents (slope undefined) is when denominator equal to zero.

OpenStudy (nincompoop):

yes

OpenStudy (perl):

so either way, you can use paul's general formula or you can do it her way. should come out the same

OpenStudy (gdgutierrez):

The answers that are shown here are Horizontal: (0,0) and (1, pi/2) Vertical: (sqrt(2)/2, pi/4) and (sqrt(2)/2, 3pi/4)

OpenStudy (gdgutierrez):

I cannot seem to know how to get that...

OpenStudy (perl):

right, so you did sin (2theta) = 0

OpenStudy (perl):

a) sin (2theta) = 0 sin^-1 (0) = 2*theta

OpenStudy (nincompoop):

by tan(x), it helps to tackle the problem intuitively because there are common graphs

OpenStudy (perl):

right, they share zero points and undefined points

OpenStudy (perl):

sin^-1(0) = 0, pi, 2pi, 3pi, ... so sin^-1(0) = 2x gives us 0,pi,2pi,3pi,...= 2x divide both sides by 2 0, pi/2, pi, 3pi/2, 2pi, 5pi/2, ...

OpenStudy (nincompoop):

i knew this polar coordinates is going to haunt me.

OpenStudy (perl):

:)

OpenStudy (perl):

so theta = 0, pi/2 (and then it repeats pi, 3pi/2, so we don't need those)

OpenStudy (gdgutierrez):

Same as vertical line right?

OpenStudy (perl):

now plug back in to find x and y x = r cos theta, x = (sin0) cos(0) y = rsin theta y = sin(0) sin(0)

OpenStudy (perl):

vertical line is when the denominator is zero , so solve cos(2theta) = 0

OpenStudy (perl):

or alternatively, when tan(2theta) is undefined

OpenStudy (perl):

So in other words, we are finding theta when dy/dx = 0 (horizontal slope) and when dy/dx is undefined (vertical slope). Once we find theta, then we can plug theta into x = r cos theta, y = r sin theta to find the x and y coordinates

OpenStudy (perl):

so theta = 0 , theta = pi/2 gives you horizontal slopes theta = pi/4 , 3pi/4 gives you vertical

OpenStudy (perl):

also we found earlier x = sin(2theta)/2 , y = sin^2(theta)

OpenStudy (perl):

horizontal -> (0,0) (0,1) vertical -> (1/2, 1/2) , (-1/2, 1/2)

OpenStudy (perl):

a fun problem is , convert this polar equation to cartesian equation, this one is doable

OpenStudy (perl):

r = sin theta multiply both sides by r r^2 = r sin theta , now we know r^2 = x^2 + y^2 and y = r sin (theta) so we have x^2 + y^2 = y x^2 + y^2 - y = 0 , and complete the square you have equation of circle